C: Short-run impacts of fiscal policy The economy of Wonderland has the following characteristics: C d = 100 + 0.5(Y − T) − 1000r I d = 400 − 1500r Md P = 0.5Y − 2000(r + π e ) π e = 0 where Y is output, T is a lump-sum tax, G is government spending, r is the real rate of interest, P is the price level. Questions 19-28 relate to this information.

C2: Fiscal policy in a small open economy with flexible exchange rate. Now suppose Wonderland trades with one large open economy which has a price level of PF or = 200. The behaviour of Wonderland can still be represented by the equations in part C1. Plus its net exports follows NX = 50 − 0.1Y − 50e where e is the real exchange rate. Questions 22-26 relate to this information
Q24. What are the implied levels of nominal exchange rate and net exports?
(a) enom = 21.1; NX = −120
(b) enom = 2; NX = −100
(c) enom = 8; NX = −180
(d) enom = 21.1; NX = −180
1
Expert's answer
2020-12-10T14:40:24-0500

Solution

C: Given the following characteristics: "C\ d = 100 + 0.5(Y − T) − 1000r\ I\ d = 400 − 1500r\ Md\ P = 0.5Y − 2000(r + π e ) π e = 0"

where

Y = output,

T = lump-sum tax,

G = government spending,

r = real rate of interest,

P is the price level.



"AD(4)=C^d+I^d+Nx+G"

AD curve; relationship between p and 4 and keep other variables as they are


"4= 100 + 0.5(Y − T) − 1000r + 400 − 1500r + 50-0.14 − 50e+G = 0 \\
4=550+0.5y-0.5T-2500r-0.1y-50e+G\\
4-0.5y+0.1y=550-0.5T-2500r-50e+G\\
0.6y=550-0.5T-2500r-50e+G\\
\therefore \frac{4^{\alpha}}{P}=0.5y-(2000)(r+r^e)\\
\therefore \pi^e=0,\\
\frac{N}{P}=0.5T-2000r\\
\implies 2000r=0.5T-\frac{M}{P}\\
\implies r=\frac{0.5T}{2000}-\frac{M}{P}\times\frac{1}{2000}\\
\therefore 0.6y=550-0.5T-2000(\frac{0.5y-\frac{M}{P}}{2000})-50e+G\\
\implies 0.6y=550-0.5T-1.25(6.5y-\frac{M}{P})-50e+G\\
\implies 0.6y=550-0.5T-0.625y+1.25\frac{M}{P})-50e+G\\
\implies 0.6y+0.625y-550-0.5T+\frac{1.25M}{P}-50e+G\\
\frac{1.25M}{P}=1.205y-550+0.5T+50e+G\\"

Multiplying (4) both sides we get


"\frac{5M}{P}=4.9y-2200+2T+200e+4G"

"p=\frac{5M}{4.94+2200+2T-4G+200e}"

For Q24, option d is correct